Suppose that, there is a perfectly competitive industry where all the firms are identical with identical cost curves. Furthermore, suppose that a representative firm’s total cost is given by the equation TC = 100 + q2 + q where q is the quantity of output produced by the firm. You also know that the market demand for this product is given by the equation P = 1000 – 2Q where Q is the market quantity. In addition you are told that the market supply curve is given by the equation

P = 100 + Q.

i. What is the equilibrium quantity and price in this market given this

Information?

ii. The firm’s MC equation based upon its TC equation is MC = 2q + 1.

Given this information, what is the firm’s profit maximizing level of production, total revenue, total cost and profit at this market equilibrium? Is this a short-run or long-run equilibrium? Explain your answer.

3. Distinguish between market period, Short-run and long run. Does the consideration of period affect the price policy?


1
Expert's answer
2021-12-21T17:11:07-0500

i) MarketEquillibrium

Ps=Pd or Qs=Qd

1000-2Q= 100+Q

Q=300

1000-2(300)= 400

P=400


ii)

MC=2q+1

Profit Maximization Output

MC=MR=P

2q+1=400

q=200.5

TR= P.Q

= 200.5*400

=80,200

TC= 1000=200.52+200.5

=41,400.75

"\Pi" =TR-TC

=80200-41400.75

=38,799.25

iii) Th firm is operating in the short run it has not yet adjusted to the market demand

3)

Short run is the period of time in which the quantity of at least one input is fixed and the quantities of the other inputs can be varied.

The long run is the period of time in which the quantities of all inputs can be varied.

The consideration of time affects price policy as at the long run quantities of all inputs have been varied unlike in the short run where one input is fixed.